OpenStudy (missykea):

Doreen Schmidt is a chemist. She needs to prepare 36 ouncesounces of a 16​% hydrochloric acid solution. Find the amount of 18​% solution and the amount of 9% solution she should mix to get this solution

OpenStudy (missykea):

please help!

OpenStudy (helpwithk12):

oreen Schmidt is a chemist. she needs to prepare 32 ounces of a 10% hydrochloric acid solution. find the amount of 16% solution and the mount of 8% solution she should mix to get this solution. I have no idea what I am doing. This is the equation I have so far. 16x+1024-32x=320

OpenStudy (missykea):

im stuck and this is my last one for the night can you help

OpenStudy (helpwithk12):

yes

OpenStudy (helpwithk12):

Let x = amount of 16% solution then 32-x = amount of 8% solution . Equation: .16x + .08(32-x) = .10(32)

OpenStudy (helpwithk12):

Do you got any answers

OpenStudy (helpwithk12):

hello

OpenStudy (amedel):

Hey if you want you could probably try to solve this one using a system of equations. Let A be the number of ounces of the 18% solution and B be the number of ounces of the 9% solution; what two facts involving A and B do we know to be true when we combine them?

OpenStudy (missykea):

im still stuck

OpenStudy (amedel):

Well we know that when we combine the two solutions their mass together will be 36 ounces right? So: A + B = 36

OpenStudy (amedel):

Additionally, if we mix it perfectly: 0.18 * A + 0.9 * B = 0.1 * 36

OpenStudy (missykea):

im still stuck

OpenStudy (amedel):

Does seeing the second equation help any?

umerlodhi (umerlodhi):

They're correct!